A lady bug is walking on the top edge of a bucket. If the radius of the opening is 8 inches and the lady bug has walked a path that is 14 inches long, what is the radian of measure of the central angle it has traveled?
Answer by stanbon(75887)

A lady bug is walking on the top edge of a bucket. If the radius of the opening is 8 inches and the lady bug has walked a path that is 14 inches long, what is the radian of measure of the central angle it has traveled?
-----------------------
circumference of the bucket is 16pi inches
------------------------------------------
Proportion EQUATIOn:
central angle/360 = 14/16pi
central angle = (14/16pi)*360 = 100.27 degrees.
====================
Cheers,
Stan H.